# DiaPer: End-to-End Neural Diarization with Perceiver-Based Attractors

### v2

@@ V-E Wide-band scenarios
Most works on end-to-end models focus on the telephone scenario and use Callhome (which is a paid dataset) as benchmark. We believe that this is partly because synthetic data (needed for training such models) match this condition quite well. However, there are many wide-band scenarios of interest when performing diarization and only few works have analyzed their systems on a wide variety of them [10, 74]. Following this direction, and pursuing a more democratic field, in this section we use DiaPer on a wide variety of corpora (most of which are of public and free access) and show the performance for the same model (before and after FT) across domains.
@@
Since most of the scenarios present many speakers per conversation, all DiaPer models were adapted to the set of 1-10 speakers per recording using sequences of 4 minutes. The 8 kHz model was trained on telephony SC and two 16 kHz models were used. Both wide-band models were trained on LibriSpeech-based SC where one model had 10 attractors (like the 8 kHz model) and another had 20 attractors to allow for more speakers. All models are evaluated without and with FT. For corpora where a multi-speaker train set is available, the train set is used for FT until no more improvements are observed on the development set. If no train set is available, the dev set is used for FT until the performance on the test set does not improve further. Therefore, results on these latter corpora should be taken with a grain of salt.
@@
Looking at the results, in some cases, there was overfitting when performing FT on the development set (since those sets did not have a train set). In DipCo, this is most likely due to the limited amount of data. In VoxConverse, the distribution of the number of speakers per recording is skewed towards more speakers in the test set and FT on the dev set makes the model find fewer speakers than without FT. Even more, recordings with more speakers are longer, making the overall error higher after FT on the test set. As for AliMeeting near mix, DiaPer (20att) has slightly worse performance on the test set but the decision to stop the FT was made observing the performance on the dev set, for which there were improvements.
@@
In comparison with the best results published at the time of writing, DiaPer performs considerably worse in most of the scenarios. However, it should be noted that in many cases the best results correspond to systems submitted to challenges which usually consist of the fusion of a few carefully tuned models. DiaPer, like any end-to-end system, is very sensitive to the type of training data. This is highly noticeable in the high errors before fine-tuning for all far-field scenarios: AISHELL-4, AliMeeting far mix, AMI mix array, CHiME6 and DipCo; and relatively lower errors for exclusively close-talk scenarios: AliMeeting near mix, AMI mix headset, Mixer6 and in the comparison between DIHARD 2 and DIHARD 3 full where the latter contains a large portion of telephone conversations. All SC (used to train the models) are generated with speech captured from short distances (telephone for the 8 kHz system and LibriSpeech for the 16 kHz ones). Using reverberation could improve the situation, but it has not been explored so far in this context. Not having enough amount of data matching the testing scenario is a strong drawback for the fine-tuning of end-to-end models as observed with DipCo and VoxConverse. Conversely, Mixer6 and RAMC with large amounts of FT data and relatively simple setups are among the scenarios with the largest relative improvement given by the FT. Even if in most cases the performance is not on par with other approaches, DiaPer’s final performance is very competitive for MSDWild and RAMC.
@@
The main goal of this comparison was to present a unified framework evaluated across different corpora. More tailored models could be trained if we used SC with specific numbers of speakers per recording (matching the evaluation data). Likewise, the output post-processing (subsampling and median filter) could be adapted for each dataset. This should definitely result in better performance and is left for future work.
@@ Acknowledgments
We thank the members of the diarization sub-group in BUT for valuable suggestions, especially Anna Silnova for feedback on the manuscript, and Dominik Klement for advice to run DVBx. We also thank Marc Delcroix, Zhengyang Chen, Shota Horiguchi, Zhihao Du, and Chin-Yi Cheng for sharing details about the number of parameters/hours of training data in some models (and all other authors for having shared that information in their work already), and Hervé Bredin for feedback on the presentation of wide-band setups.

### v3

@@ V-E Wide-band scenarios
Most works on end-to-end models focus on the telephone scenario and use Callhome (which is a paid dataset) as benchmark. We believe that this is partly because synthetic data (needed for training such models) match this condition quite well. However, there are many wide-band scenarios of interest when performing diarization and only few works have analyzed their systems on a wide variety of them [10, 74]. Following this direction, and pursuing a more democratic field, in this section we use DiaPer on a wide variety of corpora (most of which are of public and free access) and show the performance for the same model (before and after FT) across domains. The results are presented in Table XIII.
@@
Since most of the scenarios present many speakers per conversation, all DiaPer models were adapted to the set of 1-10 speakers per recording using sequences of 4 minutes. The 8 kHz model (system (2)) was trained on telephony SC and two 16 kHz models were used (systems (5) and (7)). Both wide-band models were trained on LibriSpeech-based SC where one model had 10 attractors (like the 8 kHz model) and another had 20 attractors to allow for more speakers. All models are evaluated without and with FT (systems (3), (6) and (8)). For corpora where a multi-speaker train set is available, the train set is used for FT until no more improvements are observed on the development set. If no train set is available, the dev set is used for FT until the performance on the test set does not improve further. Therefore, results on these latter corpora should be taken with a grain of salt.
@@
Looking at the results, in some cases, there was overfitting when performing FT on the development set (since those sets did not have a train set). In DipCo, this is most likely due to the limited amount of data. In VoxConverse, the distribution of the number of speakers per recording is skewed towards more speakers in the test set and FT on the dev set makes the model find fewer speakers than without FT. Even more, recordings with more speakers are longer, making the overall error higher after FT on the test set.
@@
In comparison with the best results published at the time of writing, DiaPer performs considerably worse in most of the scenarios. However, it should be noted that in many cases the best results correspond to systems submitted to challenges which usually consist of the fusion of a few carefully tuned models. DiaPer, like any end-to-end system, is very sensitive to the type of training data. This is highly noticeable in the high errors before fine-tuning for all far-field scenarios: AISHELL-4, AliMeeting far mix, AMI mix array, CHiME6 and DipCo; and relatively lower errors for exclusively close-talk scenarios: AliMeeting near mix, AMI mix headset, Mixer6 and in the comparison between DIHARD 2 and DIHARD 3 full where the latter contains a large portion of telephone conversations. All SC (used to train the models) are generated with speech captured from short distances (telephone for the 8 kHz system and LibriSpeech for the 16 kHz ones). Using reverberation could improve the situation, but it has not been explored so far in this context. Not having enough amount of data matching the testing scenario is a strong drawback for the fine-tuning of end-to-end models as observed with DipCo and VoxConverse. Conversely, Mixer6 and RAMC with large amounts of FT data (more than 100 h each) and relatively simple setups (interviews and phone calls) are among the scenarios with the largest relative improvement given by the FT. Even if in most cases the performance is not on par with other approaches, DiaPer’s final performance is very competitive for MSDWild and RAMC.
Unlike the telephony scenario where less diverse acoustic characteristics combined with plenty of data enables strong performance of EEND systems even without fine-tuning, the story with wide-band datasets is very different. This is in contrast to traditional clustering-based approaches that are quite robust to different acoustic situations present in wide-band scenarios. Current research suggests that EEND systems do not focus on learning speaker voices [92], but this might be key to make the EEND systems robust to different conditions like speaker recognition systems are. Devising models and training strategies that can bridge the gap is still an open problem. One possibility would be to capitalize on large amounts of real speaker-labeled data like those normally used to train standard clustering-based systems. Another option could be to make use of data with diarization annotations but very different characteristics.
To explore this idea, and following the strategy shown by Plaquet and Bredin [74], we pooled the sets from different corpora to generate a compound training set. Yet, as shown in Section V-D, the length of the sequences used to construct the training batches can have a relevant effect, especially in scenarios with many speakers. For this reason, three sequence lengths were explored: 1, 5 and 10 minutes corresponding to 2.7, 2.9 and 2.9 speakers per sequence on average for the whole training set. Three fine-tunings on this compound set were performed (one for each sequence length) as shown in systems (9), (10) and (11) in Table XIII. Then, for the best of the three configurations for each dataset, a final fine-tuning step was performed using only in-domain data starting from the best-performing system for that dataset among (9), (10) and (11). The sequence length used for the final fine-tuning was the same as the one corresponding to the best FT on the compound system.
In general, fine-tuning on the compound set provides gains for most sets, especially when using 5- or 10-minute sequences. This is beneficial, in particular, for DipCo and VoxConverse for which direct fine-tuning (systems (6) and (8)) does not improve the performance but FT on the compound set provides gains and even enables better performance after FT on the corresponding dev set. This strategy also provides large gains for AliMeeting far and near, AMI array and CHiME6 showing that more FT data can be beneficial on the difficult far-field scenarios.
Regarding the sequence length, it is no surprise that MSDWild reaches the best performance with 1-minute-long sequences since files are less than 1.5 minutes long on average. Surprisingly, Mixer6 also sees degradation when using longer sequences. Even more, a direct fine-tuning (system (8)) performs the best on this set. We believe this could be because there are plenty of hours of data in this relatively simple scenario, combined with the fact that the signal is obtained by combining several channels from multiple devices which could create particular conditions not seen in other datasets. Differences between (10) and (11) are in general small with AMI array and DipCo being the exceptions. While both have long recordings and far-field data which could partly explain this behavior (having longer contexts leads to better representations), the pattern is different for other datasets with similar characteristics such as AISHELL-4, AliMeeting far and, to a lesser extent, CHiME6. The impact of the FT sequence lengths was barely explored here and only some conclusions can be drawn. Nevertheless, we believe this should be better analyzed in the future to devise better training strategies adequate for a given type of data.
Handling extremely long recordings poses difficulties for DiaPer as can be seen for CHiME6. The model probably struggles to condense relevant speaker information for the whole recording on the latent space. It should be noted that the speakers in this dataset can move through different rooms and quite often they sound very quietly. DiaPer consistently finds fewer speakers than the expected four, showing that the model struggles to distinguish the voices. Mechanisms to process the input at different levels (local and global) might help addressing these issues; the clear alternative being EEND-VC-like models but modifications in the encoders to handle the input with different contexts could also provide advantages.
@@
Although we tried to shed some light on reasons for certain training strategies, we believe that many aspects need to be explored. The main goal of this comparison was to present a unified framework evaluated across different corpora. More tailored models could be trained if we used SC with specific numbers of speakers per recording (matching the evaluation data). Likewise, the output post-processing (subsampling and median filter) could be adapted for each dataset. This should definitely result in better performance and is left for future work.
@@ Acknowledgments
We thank the members of the diarization sub-group in BUT for valuable suggestions, especially Anna Silnova for feedback on the manuscript, and Dominik Klement for advice to run DVBx. We also thank Marc Delcroix, Zhengyang Chen, Shota Horiguchi, Zhihao Du, and Chin-Yi Cheng for sharing details about the number of parameters/hours of training data in some models (and all other authors for having shared that information in their work already), and Hervé Bredin for feedback on the presentation of wide-band setups. Finally, we also thank the anonymous reviewers for their comments that allowed us to improve the quality of this work.